### v1

# Measuring Fitness and Importance of Species in Food Webs

Abstract
Ecosystems face intensifying threats from climate change, overexploitation, and other human pressures, emphasizing the urgent need to identify keystone species and vulnerable ones. While established network-based measures often rely on a single metric to quantify a species’ relevance, they overlook how organisms can be both carbon providers and consumers, thus playing a dual role in food webs. Here, we introduce a novel approach that assigns each species two complementary scores—an importance measure quantifying their centrality as carbon source and a fitness measure capturing their vulnerability. We show that species with high importance are more likely to trigger co-extinctions upon removal, while high-fitness species typically endure until later stages of collapse, in line with their broader prey ranges. On the other hand, low fitness species are the most vulnerable and susceptible to extinctions. Tested on multiple food webs, our method outperforms traditional degree-based analyses and competes effectively with eigenvector-based approaches, while also providing additional insights. Relying solely on interaction data, the approach is scalable and avoids reliance on expert-driven classifications, offering a cost-effective tool for prioritizing conservation efforts.
I Introduction
In contemporary ecology, the preservation of ecosystems is an urgent and critical task. In this context, key challenges are understanding the repercussions of species loss and identifying the most vulnerable species. The significance of this task is further accentuated by the escalating anthropogenic threats, including climate change and over-exploitation of species, which strongly stress ecosystems [1, 2, 3, 4, 5, 6, 7, 8]. Species extinctions can trigger cascading effects throughout the ecosystem, generating a chain of secondary extinctions, thus amplifying the initial damage [9, 10]. To mitigate the effects of species removal, it is then crucial to quantify species “relevance” and “robustness”. This information can guide policies, focusing investments to preserve the most fundamental and less robust species.
As in many complex systems, extinctions result from non-linear complex interactions and structures [11, 12], where each component is critical for the functioning of the ecosystem. Similar behaviors are also observed, for instance, when studying supply chains, where the removal of a single firm may result in a severe output loss of the entire economy [13, 14]. In this context, the network approach has emerged as a highly effective tool to study this type of structures and to comprehend the interplay between species or, more generally, any type of interacting entities. In this representation, each species is depicted as a node, while the links between nodes symbolize their interactions, such as mutualistic, host-parasite, or trophic relationships [15, 16]. Networks thus serve as mathematical representations that map the relationships within an ecological system, encompassing various types of connections. Among the different types of networks, food webs offer a representation of “who consumes whom” interactions and have been extensively studied from the network science perspective [17, 18, 19, 20, 21, 22, 23, 24]. Network measures, especially those related to PageRank or eigenvector centrality [25, 26, 27], have demonstrated substantial efficacy in quantifying the relevance of species in food webs [28, 29], identifying those whose removal would trigger the most significant impact on overall biodiversity. In particular, network measures prove particularly invaluable when considering large food webs, where exhaustive consideration of all potential extinction sequences is computationally prohibitive. However, while these approaches help detect crucial species, they fail to identify the most vulnerable ones. Indeed, many species play little or no role in the food web economy, but could still be in severe danger. This is, for instance, the case for pandas. Their extinction would not cause a direct secondary cascade in the food web, but their very non-diverse diet makes them extremely susceptible to changes in the ecosystem.
In light of these considerations, our research introduces a novel methodology for measuring a species’ dual nature in food webs: both as carbon source and consumer. Indeed most species absorb carbon from other species but are also consumed by larger predators or, eventually, by decomposers. This methodology quantifies the importance (susceptibility to predation and relevance for the food web) and fitness (predatory prowess and robustness to extinctions) of each species. By employing these two quantitative measures, we move beyond traditional one-dimensional evaluations that fail to capture the complex dual role of species in food webs. Conversely, our approach provides a holistic visualization of the entire food web, encapsulating its complex structure within a bidimensional plane, where each species is represented as a point. Our analysis demonstrates that important species are also those leading to larger extinction cascades, while low-fitness species possess little resistance to such events. Our methodology, therefore, enables a more accurate identification of both critical and vulnerable species within ecosystems, with the potential of greatly enhancing the effectiveness of conservation strategies.
Figure 1: Cypress food web: network representations and Fitness-Importance plane. a, Directed monopartite subnetwork illustrating predator-prey relationships. b, Equivalent undirected bipartite representation of the subnetwork. c, Fitness-Importance plane depicting species’ ecological roles. Points represent individual species, with colors indicating functional categories and sizes proportional to species biomass. Selected species are labeled for reference. This visualization reveals the dual roles of species as both predators and prey within the ecosystem.
II Results
II.1 The Fitness and Importance of Species
Quantifying species’ roles and impacts within ecosystems presents a significant methodological challenge in ecological research. Traditional network-based methodologies approach this task by looking at the food web formed by the species. This consists of a directed network, where nodes are the species, while edges signify carbon exchange between these species. Being predation one of the main carbon exchange mechanisms, these graphs are often also called prey-predator networks. We show in Fig. 1a a schematized food web, where we adopt the convention of arrows going from predators to prey. Once a food web has been reconstructed, the standard approach consists in applying a centrality measure to rank the species. For instance, eigenvector centrality has proven very effective in determining which species have a stronger potential to cause chain extinctions if removed from an ecosystem [28]. These measures, however, are unable to assess which species are most in danger during these extinction cascades.
To address this limitation, we introduce a novel algorithmic approach that quantifies species’ dual roles through two distinct measures: importance and fitness. Importance measures how important in the food web a species is, while fitness assesses a species’ predatory capabilities and adaptability within the food web. The importance of a species is high if it serves as prey for multiple species with low fitness scores. On the other hand, fitness evaluates a species’ predatory efficiency by assessing both the quantity and importance of its prey species. Species that prey upon diverse taxa, particularly those of low importance, demonstrate high predatory capacity and consequently receive high fitness values. This bi-dimensional characterization enables a more comprehensive understanding of species’ ecological roles compared to traditional network-based metrics.
Mathematically, we formulate this framework as a bi-dimensional non-linear map. Let $\bm{M}$ represent the adjacency matrix of the food web. The element $M_{ij}$ equals one if there is carbon transfer (predation) from species $i$ to $j$, and zero otherwise. The fitness ($F_{i}$) and importance ($I_{i}$) of species $i$ are computed through the following iterative map:
$\left\{\begin{array}[]{lcl}F_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ji}\,/\,{I_{j}^{(n)}}\\ I_{i}^{(n+1)}&=&\delta+\sum_{j}M_{ij}\,/\,{F_{j}^{(n)}}\end{array}\right.F_{i}^{(0)}=I_{i}^{(0)}=1\leavevmode\nobreak\ \forall i.$ (1)
Here, $F_{i}^{(n)}$ and $I_{i}^{(n)}$ denote the fitness and importance of species $i$ at iteration $n$, respectively. The quantity $\delta$ is a regularization term that guarantees convergence. It does not affect the final ranking, as soon as it is way smaller than $\mathbf{M}$’s elements. In this paper, we set it at $10^{-3}$. The algorithm iterates until the convergence of both quantities. The first equation is telling us that a species has high fitness if it can absorb carbon by several sources, in particular from those with low importance. The latter are indeed only consumed by a small set of other species and are thus expected to be harder prey. The second equation implies that the importance of a species is high if it is consumed by several other species, in particular those with low fitness. Low fitness species can indeed only absorb carbon from a limited set of sources, so the extinction of any of those would put them in serious danger. For this reason, we identify vulnerability as the inverse of fitness. The lower the fitness of a species, the more vulnerable to food web shocks it is expected to be.
This approach shares notable similarities with the Economic Fitness and Complexity (EFC) algorithm developed for bipartite networks [30, 31, 32, 33]. While the EFC algorithm has successfully analyzed various bipartite nested networks, including ecological systems [16], its application to mono-partite directed networks represents a novel extension [34]. In ecological contexts, nestedness contributes significantly to ecosystem stability, biodiversity, and resilience, particularly in mutualistic networks where specialist species typically interact with generalist ones [35]. This similarity between the two approaches is not surprising, since, as shown in Fig. 1b, a food web can also be represented as a bipartite network. Differently from standard bipartite networks, layers contain the same set of species. The first looks at their role as predator, and the second one looks at their role as prey. Links between the two layers connect predators to their prey.
II.2 Fitness-Importance Plane
Figure 2: Fitness-Importance plane. In royal blue, the marine ecosystems; in khaki, the terrestrial one; in medium aquamarine the mixed ones. Values are log-transformed and normalized to [0,1] for cross-food web comparison. The plane is divided into four quadrants. Base Trophic Nodes (top left quadrant), Carbon Bridge Nodes (top right quadrant), Sheltered Trophic Nodes (bottom left quadrant), Apex Trophic Nodes (bottom right quadrant).
Our algorithm allows us to place all the species within an ecosystem on the 2-dimensional space defined by the fitness and the importance. Each axis captures a different aspect of species, namely their importance as prey and carbon suppliers and their prowess as predators and carbon absorbers. As an example, we show in Fig. 1c the fitness-importance plane for the Cypress dry season food web [36]. In this figure, each point corresponds to a species, different colors/shapes correspond to five different categories of species (see Methods for details), while point size is proportional to the biomass. Remarkably, species are not randomly scattered on the plane but tend to cluster following the classification. Producers and Organic Matter (dark green circles) are mostly placed on the top left corner since they are crucial carbon suppliers (high importance) and they do not absorb carbon from other species (low fitness). Conversely, large predators are placed in the bottom right corner, since they are both very hard to prey on and very good at preying. Also, a large fraction of secondary consumers are located in this region. The top right corner contains species that can absorb carbon from several other species, but that, at the same time, also transfer carbon to many other species. Not surprisingly in this area we find invertebrates and detritivores or mammals like rats. Finally, the bottom left corner mostly contains herbivores/omnivores and secondary consumers. These are species such as lizards that are not central in the food chain functioning, being hard to prey on, but with little to none preying capabilities.
Food web In-degree (%) Eigenvector-in (%) Importance (%)
St. Marks 84.8 $\pm$ 1.1 91.6 91.4
Benguela 77.3 $\pm$ 0.5 94.6 91.3
Coachella 70.9 $\pm$ 0.3 75.9 78.8
Florida 64.6 $\pm$ 0.3 82.4 82.6
Cypress 78.5 $\pm$ 0.4 81.9 86.9
Mangrove 72.5 $\pm$ 0.1 95.0 94.3
Table 1: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance rankings at the outset and keep them fixed throughout the simulated extinction process. Values represent means and standard deviation. For eigenvector-in and importance, the standard deviation is smaller than the last significative digit displayed. Higher percentages indicate greater ecosystem sensitivity to species loss based on the respective ranking method. Bold values indicate the best-performing method for each ecosystem.
As we show in Fig. 2, we can then approximately divide the fitness-importance plane into four regions:
  • Base Trophic Nodes (top left quadrant)
    These species are characterized by poor predatory ability but are heavily preyed upon by a multitude of species. They form the foundation for energy and nutrient flow, providing a primary source of food for consumers. They are both very important and very vulnerable (low fitness).
  • Carbon Bridge Nodes (top right quadrant)
    These species excel at absorbing or transforming carbon (often from detrital or non-living sources) and are widely consumed by other species. They “bridge” carbon from lower-level or non-living pools into higher trophic levels. Also them are very importance, but not as vulnerable, since their high fitness allows them to absorb carbon from very diverse sources.
  • Sheltered Trophic Nodes (bottom left quadrant)
    These species are neither strong predators nor frequent prey; they have defenses, behaviors, or ecological niches that protect them from heavy predation. They have low fitness, so high vulnerability, but their extinction would likely not cause cascades.
  • Apex Trophic Nodes (bottom right quadrant)
    These species possess high predatory ability with low vulnerability. They regulate prey populations and structure ecosystem dynamics from the top of the food web. They have low importance, but also low vulnerability.
Fig. 2 is realized considering six different ecosystems from terrestrial, aquatic, and mixed domains. Note that we normalized fitness and importance to compare species from different ecosystems (details in Methods). Colors denote the ecosystem type, while symbols show the different food webs. We observe a predominance of species in the Base Trophic Nodes ($42.09\%$) and Apex Trophic Nodes ($32.65\%$) quadrants, only a few Carbon Bridge Nodes ($2.30\%$) and a minority of Sheltered Tropic Nodes ($22.96\%$). We report in SI, for each ecosystem, the detailed list of all species in each quadrant. It is important to stress that fitness and importance only present a marginal anti-correlation, confirming that the two quantities actually capture complementary species’ properties.
II.3 Cascading Extinctions and Robustness of Species
Figure 3: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log fitness of extinct species vs. extinction step, based on importance ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log fitness vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison. This plot illustrates ecosystem resilience patterns and differential impacts of species loss across varied food web structures.
The picture suggested by the fitness-importance plane, while reasonable, needs stronger validation. In particular, we need to understand whether the two measures we introduced are correlated with the predatory capabilities of animals and their vulnerability and importance as prey. These are not easy quantities to measure, but we can indirectly assess them.
We start studying importance, which, as mentioned, is a measure of a species’ relevance as a carbon source. High-importance species play a major role in food webs since they represent the base of the food chain. This role can be quantified looking at the so called extinction curve of a food web. Such a curve is obtained by progressively removing species from the food web and looking at which other species get disconnected, thus remaining without any supplier. More details are reported in the Methods section. The premise is that a more accurate measure of species relevance within a food web should lead to a faster network collapse when species are removed in order of their ranking. Indeed, Eq. (1) implies that a species with high importance is preyed on by several low-fitness species. Therefore, its removal would likely cause many of these low-fitness, low-predatory capabilities species to be unable to sustain themselves. We quantify this effect using the mean extinction area (area under the extinction curve), where higher values indicate a better ability to identify critical species within the ecosystem. Table 1 presents the mean extinction areas following 3 different ranking strategies: in-degree, eigenvector-in centrality, and importance. We establish the initial rankings and keep them fixed throughout the simulated extinction process, whereas in the Supplementary Information, we present the scenario where rankings are recalculated after each node removal, including extinction curves plots for both cases across all food webs. The results reveal that both eigenvector-in and importance consistently outperform in-degree centrality across all studied ecosystems. Notably, eigenvector-in centrality and importance demonstrate comparable efficacy, with each measure excelling in different ecosystems. These results validate importance, confirming its ability to identify the most crucial species within food webs.
As a second step, we study the ability of fitness to measure the vulnerability of species. A high-fitness species can prey on many other organisms, including those with low importance, i.e., those that are very hard to prey on. As a consequence, we expect high-fitness species to be less affected by extinctions within the ecosystem. They are indeed very diversified and less susceptible to such events. On the other hand, low-fitness species are expected to be the most vulnerable and thus the first to get extinct during cascade events. We can measure this by looking at the average (log) fitness of species that get extinct at a given step of the extinction curve construction. As shown in Fig. 3a for the cypress food web, this average increases with the step, meaning that high-fitness species only get extinct when most other species are already gone. We report next to each point a sample species going extinct in that step, while the inset shows the same plot, but using the average eigenvector-out centrality to measure (inverse) vulnerability. A much weaker correlation is observed. We repeat the same analysis for all the six food webs we considered, reporting the results in Fig. 3b. Note that we normalized both the step and the fitness for comparison reasons. In all cases, we observe the same behavior. More quantitatively, we can compute the Spearman correlation between the removal step and the (average log) fitness of the removed species. The average of this quantity among all ecosystems is $0.98$, while the average of the eigenvector-out centrality is $0.72$, thus consistently lower. We repeat the analysis using eigenvector-in centrality for species removal, yielding similar results that support fitness as an effective measure of species’ predatory skills and extinction resilience. In the Supplementary Information, we report details on Spearman correlation coefficients and eigenvector-in-based removals, along with additional comparisons between fitness, importance, and other centrality measures.
III Discussion
Ecosystems face intensifying threats, making it urgent to identify vulnerable species, to prevent their extinction. Current network-based approaches often reduce the complex predator-prey interactions to a single metric. While this allows us to understand which species are crucial for the functioning of a food web, it tells little to nothing about the vulnerability of species. Our method, based on a bi-dimensional non-linear map, addresses this gap, associating to each species both a importance and a fitness. The former is linked to the susceptibility to predation and measures how crucial a species is in the food web. The latter measures predatory prowess and quantifies the robustness of species. High-importance species, which often anchor lower trophic levels, appear critical to maintaining overall network stability, and their loss disproportionately impacts a wide range of other species. On the other hand, high-fitness species thrive on diverse prey pools, making them more resistant to ecosystem perturbations. Our analysis confirms this intuition, showing that importance effectively identifies species whose removal triggers significant co-extinctions. At the same time, fitness highlights species capable of persisting until the late stages of collapse, with the most vulnerable, low-fitness, species being the first to disappear. While eigenvector-in centrality can resemble our importance ranking, neither eigenvector-in nor eigenvector-out centralities adequately capture the predatory strength and resilience that define a species’ fitness. This limitation arises from the linear nature of the eigenvector centrality, which is often incapable of capturing complex, non-linear interactions.
By mapping species onto a fitness-importance plane, we can reveal patterns that align with known functional roles (e.g., apex predators, base trophic nodes). While standard approaches rely on classifications by experts, our methodology is entirely algorithmic and data-driven. This framework holds clear promise for conservation and ecosystem management, where limited resources demand strategic prioritization. By jointly quantifying species’ importance and fitness, practitioners can pinpoint “keystone” nodes —those whose loss risks triggering wider co-extinctions— thereby directing scarce funding and protective measures where they have the greatest impact. Particular attention can also be devoted to those species that despite a marginal role as carbon suppliers, are in severe danger due to their low fitness. Because our algorithmic approach draws solely on quantitative interaction data, it avoids reliance on expert-based classifications and readily scales to large, complex food webs. When combined with objective risk assessment, this scalability ensures broader applicability across diverse ecosystems, ultimately supporting more targeted and cost-effective biodiversity preservation efforts.
IV Methods
IV.1 Data
We analyzed the following seven food webs:
Florida Bay
This network represents trophic flows through the ecosystem of Florida Bay. It includes various species such as seagrass and algae producers, microfauna, macroinvertebrates, fishes, mammals, and avifauna [36].
Cypress Dry Season
This network depicts the cypress wetlands in South Florida during the dry season [36].
Mangrove Wet Season
This network illustrates the mangrove ecosystem in South Florida during the wet season [36].
St. Marks
This network describes a winter’s Halodule wrightii (shoal grass) community in Goose Creek Bay, St. Marks National Wildlife Refuge, Florida [37].
Halodule wrightii
(shoal grass) community in Goose Creek Bay, St. Marks National Wildlife Refuge, Florida [37].
Coachella
This network represents the sand community in the Coachella Valley desert. The biota includes species of vascular plants, vertebrates, arachnids, microorganisms, and insects [38].
Benguela
This network characterizes the Benguela Marine Ecosystem, including marine mammals and fisheries [39].
IV.2 Networks Construction and Preprocessing
Food webs are represented as directed monopartite networks, where edges denote energy flow between species. In our binary model, an edge from node $i$ to node $j$ indicates that species $i$ preys upon species $j$. Self-loops are included to account for cannibalism within species. We selected six food webs for analysis and modify them to ensure irreducibility and primitivity, following the method described by [28, 40, 41]. This modification process involved:
  1. 1.
    Adding a root node to represent the external environment.
  2. 2.
    Creating links from primary producers to the root, reflecting the origin of matter in the food web.
  3. 3.
    Establishing links from the root to every node, representing intrinsic matter loss and detritus accumulation, which is partially recycled back into the food web.
To maintain consistency across datasets, we remove three nodes (Input, Output, and Respiration) from the Florida Bay, Cypress Dry Season, and Mangrove Wet Season networks. This preprocessing step ensures that all networks represent comparable ecological entities. The macroscopic properties of these modified networks, including the number of nodes, number of links, and network density, are detailed in the Supplementary Information.
IV.3 Fitness-Importance Plane
To categorize the species of the cypress food web into 5 categories, we use ChatGPT-4o (accessed on 24/10/2024). In particular, we use the following prompt, Can you classify the nodes in this food web into 5 categories?, followed by the list of species in the Cypress food web. The SI provides complete species categorizations.
To enable cross-network comparisons, we normalize the fitness and importance values for each food web using the following procedure:
  1. 1.
    Compute fitness and importance for each species.
  2. 2.
    Take the logarithm of these values.
  3. 3.
    Scale the log-transformed values to the [0,1] interval using min-max normalization:
    $x_{\mathrm{normalized}}=\frac{\log(x)-\min(\log(x))}{\max(\log(x))-\min(\log(x))}.$
This approach preserves relative differences while allowing for consistent comparison across different food webs.
IV.4 Extinction Areas
We implement the following species removal procedure to assess ecosystem robustness. Species are progressively removed based on rankings from specified algorithms, maintaining fixed rankings throughout the process. Cascading extinctions are monitored, with a species considered extinct when left without outgoing links or with only a self-loop. The procedure continues until all species go extinct. An extinction curve is generated by plotting the fraction of extinct species against the number of removed species. The extinction area, calculated as the integral of this curve, measures the ecosystem vulnerability. This area equals 1 when all species go extinct after the first removal and approaches 0.5 when no secondary extinctions occur. To address ranking degeneracies, we average results from 200 randomized tie-breakers and compare rankings based on importance, eigenvector-in, and in-degree centralities. For eigenvector-based rankings, we follow the method in [28]: at each step, we compute the dominant eigenvector of the column-normalized adjacency matrix and remove the highest-scoring node. The dominant eigenvalue is always 1, with its associated positive eigenvector summing to 1. For eigenvector-out rankings, we use the transpose of the adjacency matrix.
Supplementary Information
Appendix S1 Network Statistics
Table 2 presents key statistics for each food web, including the number of nodes ($n$), number of links ($m$), and network density ($d$). The number of nodes includes an additional root node for each network, and the number of links considers self-loops. Network density is calculated as $d=\frac{m}{n^{2}}$ for directed networks with self-loops.
Food Web # nodes # links density (%)
St. Marks 49 275 11.5
Benguela 30 234 26.0
Coachella 30 294 32.7
Florida 126 2078 13.1
Cypress 69 634 13.3
Mangrove 95 1439 15.9
Table 2: Network statistics for analyzed food webs.
Appendix S2 Cypress Food Web: Chat GPT Categories
The following are the five categories into which the species of the Cypress food web have been divided, as generated by Chat GPT.
  • Producers/Organic Matter: Living POC, Living sediment, Phytoplankton, Float. vegetation, Periphyton/Macroalgae, Macrophytes, Epiphytes, Understory, Vine Leaves, Hardwoods Leaves, Cypress Leaves, Cypress Wood, HW Wood, Roots
  • Invertebrates/Detritivores: Crayfish, Apple Snail, Prawn, Aquatic Invertebrates, Ter. Invertebrates, Refractory Det., Liable Det., Vertebrate Det.
  • Herbivores/Omnivores: Small Fish, herb and omniv, Small Frogs, Tadpoles, Galliformes, White ibis, Gruiformes, Caprimulgiformes, Hummingbirds, Woodpeckers, Passeriformes onniv., Squirrels, Mice & Rats, Rabbits, White-Tailed Deer, Armadillo
  • Secondary Consumers: Small Fish, prim. carniv, Large Fish, Turtles, Lizards, Snakes, Salamanders, Large Frogs, Medium Frogs, Salamander L, Pelecaniformes, Anseriformes, Vultures, Egrets, Great blue heron, Other herons, Wood stork, Passeriformes pred., Opossum, Shrews, Bats, Mink, Otter, Bobcat
  • Large Predators: Alligators, Kites & Hawks, Owls, Black Bear, G. Fox, Raccoon, Florida Panther, Hogs
Appendix S3 Fitness and Importance
S3.1 Quadrants Fitness-Importance Plane
Species in each food web are categorized into four quadrants based on their fitness and importance: Base Trophic Nodes, Sheltered Trophic Nodes, Apex Trophic Nodes, and Carbon Bridge Nodes. The full species lists for each category are provided in Appendix S6.
S3.2 Correlation with Eigenvector Centralities
Figure 4 reveals a strong correlation between importance and eigenvector-in centrality (Spearman coefficient = 0.93), explaining their similar performance in extinction area analysis. Conversely, Figure 5 shows a weak correlation between fitness and eigenvector-out centrality (Spearman coefficient = 0.22), accounting for significant differences in identifying resilient species.
Figure 4: Relationship between eigenvector-in centrality and importance in the Cypress food web. The Spearman correlation coefficient (s) is reported.
Figure 5: Relationship between eigenvector-out centrality and fitness in the Cypress food web. The Spearman correlation coefficient (s) is reported.
Appendix S4 Cascading Extinctions
S4.1 Recomputed Rankings
Table 3 compares mean extinction areas using different ranking strategies, recomputing rankings after each species removal. The results reveal that in this case eigenvector-in centrality generally outperforms importance, which in turn outperforms in-degree. Notably, for the Coachella food web, importance shows superior performance compared to eigenvector-in centrality.
Food Web In-degree (%) Eigenvector-in (%) Importance (%)
St. Marks 84.2 $\pm$ 1.1 91.5 87.4
Benguela 78.2 $\pm$ 0.3 94.7 87.9
Coachella 71.5 $\pm$ 0.1 71.6 79.3
Florida 63.5 $\pm$ 0.0 87.3 77.8
Cypress 78.6 $\pm$ 0.3 86.6 82.2
Mangrove 68.1 $\pm$ 0.4 95.3 92.3
Table 3: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance rankings at the outset and recompute them after each step of the simulated extinction process. Values represent means and standard deviation. For eigenvector-in and importance, the standard deviation is smaller than the last significative digit displayed (except for the eigenvector-in in the Coachella food web, for which it is 1.0). Bold values indicate the best-performing method for each ecosystem.
Food Web In-degree (%) Eigenvector-in (%) Importance (%)
St. Marks 84.8 $\pm$ 1.1 91.6 91.4
Benguela 78.2 $\pm$ 0.3 94.7 91.3
Coachella 71.5 $\pm$ 0.1 75.9 79.3
Florida 64.6 $\pm$ 0.3 87.3 82.6
Cypress 78.6 $\pm$ 0.3 86.6 86.9
Mangrove 72.5 $\pm$ 0.1 95.3 94.3
Table 4: Comparison of maximum mean extinction areas across ecosystems. For each food web and ranking method (in-degree, eigenvector-in, and importance), the table presents the higher mean extinction area obtained from either the fixed or recomputed ranking strategy.
S4.2 Recomputed vs. Fixed Rankings
Table 4 shows the maximum mean extinction areas for each food web and ranking method (in-degree, eigenvector-in, importance). The reported value represents the better result between the fixed and recomputed ranking strategies.
In Figure 6, we compare the extinction curves for all food webs under two removal strategies (recomputed and fixed initial ranking) for importance ranking. Coachella is the only food web where recomputing outperforms the fixed rankings.
Figure 6: Comparative extinction curves across food webs. Extinction trajectories for all six food webs, illustrating species loss under two removal strategies: dynamically recomputed importance ranking and fixed initial importance ranking. Each subplot depicts the proportion of extinct species versus the proportion of species removed.
Appendix S5 Robustness of Species
Table 5 presents the Spearman rank correlation coefficients between removal steps and two ecosystem metrics: average log fitness and average log eigenvector-out centrality of extinct species. These coefficients were calculated based on species removals following the importance ranking. The results demonstrate that fitness consistently outperforms eigenvector-out centrality in identifying resilient species across all food webs.
Food Web Fitness (%) Eigenvector-out (%)
St. Marks 100.00 92.86
Benguela 100.00 80.00
Coachella 100.00 60.71
Florida 91.93 55.09
Cypress 99.09 54.55
Mangrove 100.00 90.48
Table 5: Spearman rank correlation coefficients (%) between removal steps and two ecosystem metrics: average log fitness and average log eigenvector-out of extinct species. Removals were performed following the importance ranking.
This superior performance of fitness holds true even when species removals are based on eigenvector-in centrality, as illustrated in Figure 7a for the Cypress food web. Figure 7b extends this analysis to all six food webs, showcasing normalized average log fitness versus normalized extinction steps. This comparison reveals consistent patterns of ecosystem resilience across diverse food web structures.
Figure 7: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log fitness of extinct species vs. extinction step, based on eigenvector-in ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log fitness vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison.
In the specific case of the Cypress food web, Figure 8 provides additional insights into the extinction dynamics. The visualization reveals that large predators, primarily Apex Trophic Nodes, persist until later removal stages, consistent with their high fitness scores and broad prey bases.
Figure 8: Cumulative species extinctions by functional category in the Cypress food web, based on importance ranking. Stacked bar chart illustrating the progressive accumulation of extinct species at each removal step. Colors represent different functional categories. Bar heights remain constant when no new extinctions occur, reflecting the cumulative nature of the data. This visualization demonstrates the differential impact of species removals on various functional groups within the ecosystem.
Appendix S6 Fitness-Importance Species Categorization
Tables 6 and 7 present the species categorization for each food web based on their position in the fitness-importance plane. The Carbon Bridge Nodes for each relevant food web are as follows:
  • Florida Bay: Benthic POC and Water POC.
  • Cypress: Terrestrial Invertebrates, Liable Detritus, Refractory Detritus, and Vertebrate Detritus.
  • Mangrove: POC, INSCT, and C in SED.
Food Web Base Trophic Nodes Sheltered Trophic Nodes Apex Trophic Nodes St. Marks Benthic bacteria, Microfauna, Meiofauna, Bacterioplankton, Microprotozoa, Epiphyte-grazing amphipods, Suspension-feeding molluscs, Deposit-feeding peracaridan crustacean, Deposit-feeding gastropods, Other gastropods, Deposit-feeding polychaetes, Zooplankton, Halodule, Micro-epiphytes, Macro-epiphytes, Benthic algae, Phytoplankton, Detritus Hermit crabs, Spider crabs (herbivores), Isopods, Brittle stars, Herbivorous shrimps, Tonguefish, Sheepshead minnow, Epiphyte-grazing gastropods, Suspension-feeding polychaetes Omnivorous crabs, Blue crabs, Predatory shrimps, Catfish and stingrays, Gulf flounder and needlefish, Southern hake and sea robins, Atlantic silverside and bay anchovies, Killifishes, Gobies and blennies, Pipefish and seahorses, Red drum, Predatory gastropods, Predatory polychaetes, Benthos-eating birds, Fish-eating birds, Fish and crustacean-eating birds, Gulls, Raptors, Herbivorous ducks Benguela Phytoplankton, Benthic filter-feeders, Bacteria, Microzooplankton, Mesozooplankton, Macrozooplankton Benthic carnivores, Gelatinous zooplankton, Anchovy, Pilchard, Round herring, Lightfish, Lanternfish, Goby Other pelagics, Horse mackerel, Chub mackerel, Other groundfish, Hakes, Squid, Tunas, Snoek, Kob, Yellowtail, Geelbek, Whales & dolphins, Birds, Seals, Sharks Coachella plants/plant products, detritus, carrion, soil microbes soil microarthropods and nematodes, surface arthropod detritivores, surface arthropod herbivores, herbivorous mammals and reptiles soil micropredators, soil macroarthropods, soil macroarthropod predators, small arthropod predators, medium arthropod predators, facultative arthropod predators, life-history arthropod omnivore, spider parasitoids, primary parasitoids, hyperparasitoids, facultative hyperparasitoids, primarily herbivorous mammals and birds, small omnivorous mammals and birds, predaceous mammals and birds, arthropodivorous snakes, primarily arthropodivorous lizards, primarily carnivorous lizards, primarily carnivorous snakes, large primarily predaceous birds, large primarily predaceous mammals, golden eagle
Table 6: Species categorization for St. Marks, Benguela, and Coachella food webs based on their trophic roles.
Food Web Base Trophic Nodes Sheltered Trophic Nodes Apex Trophic Nodes Florida Bay $2\mu$m Spherical Phytoplankt, Synedococcus, Oscillatoria, Small Diatoms ($<20\mu$m), Big Diatoms ($>20\mu$m), Dinoflagellates, Other Phytoplankton, Benthic Phytoplankton, Thalassia, Halodule, Syringodium, Drift Algae, Epiphytes, Water Flagellates, Water Cilitaes, Acartia Tonsa, Oithona nana, Paracalanus, Other Copepoda, Meroplankton, Other Zooplankton, Benthic Flagellates, Benthic Ciliates, Meiofauna, Sponges, Coral, Other Cnidaridae, Echinoderma, Bivalves, Detritivorous Gastropods, Epiphytic Gastropods, Predatory Gastropods, Detritivorous Polychaetes, Predatory Polychaetes, Suspension Feeding Polych, Macrobenthos, Benthic Crustaceans, Detritivorous Amphipods, Herbivorous Amphipods, Isopods, Herbivorous Shrimp, Predatory Shrimp, Pink Shrimp, Thor Floridanus, Lobster, Detritivorous Crabs, Omnivorous Crabs, Predatory Crabs, Callinectus sapidus, Stone Crab, Sardines, Anchovy, Bay Anchovy, Lizardfish, Catfish, Eels, Toadfish, Brotalus, Halfbeaks, Needlefish, Other Killifish, Goldspotted killifish, Rainwater killifish, Sailfin Molly, Silverside, Other Horsefish, Gulf Pipefish, Dwarf Seahorse, Other Snapper, Mojarra, Grunt, Pinfish, Scianids, Parrotfish, Mullet, Blennies, Code Goby, Clown Goby, Flatfish, Filefishes, Puffer, Other Pelagic Fishes, Other Demersal Fishes Roots, Sharks, Rays, Bonefish, Snook, Grouper, Jacks, Pompano, Gray Snapper, Porgy, Spotted Seatrout, Red Drum, Spadefish, Mackerel, Big Herons & Egrets, Small Herons & Egrets, Ibis, Roseate Spoonbill, Herbivorous Ducks, Omnivorous Ducks, Gruiformes, Small Shorebirds, Gulls & Terns, Kingfisher, Loggerhead Turtle, Green Turtle, Hawksbill Turtle, Manatee, DOC Tarpon, Barracuda, Loon, Greeb, Pelican, Comorant, Predatory Ducks, Raptors, Crocodiles, Dolphin Cypress Living POC, Living sediment, Phytoplankton, Float. vegetation, Periphyton/Macroalgae, Macrophytes, Epiphytes, Understory, Hardwoods Leaves, Crayfish, Apple Snail, Prawn, Aquatic Invertebrates, Small Fish herb and omniv, Passeriformes onniv. Vine Leaves, Cypress Leaves, Cypress Wood, HW Wood, Roots, Lizards, Medium Frogs, Small Frogs, Salamander L, Tadpoles, Anseriformes, Vultures, Galliformes, Caprimulgiformes, Hummingbirds, Woodpeckers, Passeriformes pred., Opossum, Shrews, Bats, Rabbits, White-Tailed Deer Small Fish prim. carniv, Large Fish, Alligators, Turtles, Snakes, Salamanders, Large Frogs, Pelecaniformes, Kites & Hawks, Egrets, Great blue heron, Other herons, Wood stork, White ibis, Gruiformes, Owls, Black Bear, G. Fox, Raccoon, Mink, Otter, Florida Panther, Bobcat, Squirrels, Hogs, Armadillo Mangrove PHY, OTH. PP, LEAF, MICR. HO, ZOOPL., BACT.SED., FLA. SED., CIL. SED., MEIOF., MERO, EPIFN, POLY, AGAST, BVLVS, MBENTH, SCRUST, AMPHI, PENAID, CARID, OSHMP, JLOBST, OCRAB, TCRAB, DCRAB, PCRAB, FWINV, LARV, SPIDR, HERR, ANCH, KILLI, POEC, HRSE, SLVR, FWFSH, MOJA, GOBY, DOC WOOD, ROOT, TGAST, RAYS, EFISH, PIN, MULL, LZRD, VULT, C & C, WOODP, PASSOMN, PASSPERD, RABT, SQIUR, M & R, DERS, MANA SHRK, TARP, NEED, SNOOK, CUDA, SNAP, SCIAE, OFISH, TURT, SNKS, COCO, AMPH, L & G, PELC, CORM, BH & E, SE & E, IBIS, DUCK, DUCK, DUCK, K & H, MRAPT, GUIF, SSBIRDS, G & T, OWLS, OPSU, FOX, BEAR, RACO, M & O, CATS, DOLP
Table 7: Species categorization for Florida Bay, Cypress, and Mangrove food webs based on their trophic roles.